# HunyuanProver: A Scalable Data Synthesis Framework and Guided Tree Search for Automated Theorem Proving

### v1

@@ Abstract
We introduce HunyuanProver, an language model finetuned from the Hunyuan 7B for interactive automatic theorem proving with LEAN4. To alleviate the data sparsity issue, we design a scalable framework to iterative synthesize data with low cost. Besides, guided tree search algorithms are designed to enable effective “system 2 thinking“ of the prover. HunyuanProver achieves state-of-the-art (SOTA) performances on major benchmarks. Specifically, it achieves a pass of 68.4% on the miniF2F-test compared to 65.9%, the current SOTA results. It proves 4 IMO statements (imo_1960_p2, imo_1962_p2, imo_1964_p2 and imo_1983_p6) in miniF2F-test. To benefit the community, we will open-source a dataset of 30k synthesized instances, where each instance contains the original question in natural language, the converted statement by autoformalization, and the proof by HunyuanProver.
@@ 1 Introduction
We propose HunyuanProver, a framework designed to address the aforementioned challenges in automatic theorem proving. HunyuanProver takes two core modules: a scalable prover-data generator and guided tree-search algorithms. The prover-data generator only takes open-source data to train the initial autoformalizer and prover. The autoformalizer then converts a large amount of existing math questions into the format of the target prover (e.g., LEAN4). The prover is then iteratively improved on such data where new proof data is generated at each iteration to train the prover. At testing time, tree-search algorithms and multiple critic models are designed to perform “slow thinking” that is empirically essential for solving complex theorem-proving tasks.
@@
  • We introduce a scalable pipeline that utilizes open-source prover data and massive math problems in natural language to generate a large number of new training data for automatic theorem proving.
@@
  • We will open-source approximately 30,000 data instances at XXX. Each instance includes the formal statement, proofs searched by HunyuanProver, and the original problems.
@@ 4.1 Setup
We use the LEAN engine from LeanDojo (Yang et al., 2024) as the engine to conduct both tactic data generation and benchmark evaluation. Whole and per-step timeout for tactic executing in LEAN are set to 3600 seconds and 60 seconds, respectively. At most 800 search steps are conducted for both BFS and MCTS. For each search step, $K=8$ tactics are sampled from the given LEAN state under temperatures 0.7, 0.8, 1.0 and 1.1, where two tactic are sampled under each temperature. These temperature values are empirically decided.
@@ 4.2 Results and Analysis
Table 1 shows the comparison between HunyuanProver with other existing state-of-the-art systems. Among these systems, our HunyuanProver with distance-tritic as guidance shows the best accuracy on miniF2F, advancing the previous SOTA system (InternLM2.5-StepProver+BFS+CG) by 2.5% points using less search budget. Though InternLM2.5-StepProver+BFS+CG is slightly better than DeepSeek-Prover-V1.5-RL+MCTS, the later does not take any critic for guidance. We believe MCTS can be more beneficial to the task of automatic theorem proving, and this potential has not been fully revealed. In addition, we can also probe the importance of critic guidance by incorporating DCG on HunyuanProver, which shows 3.6%-point gain.

### v3

@@ Abstract
We introduce HunyuanProver, an language model finetuned from the Hunyuan 7B for interactive automatic theorem proving with LEAN4. To alleviate the data sparsity issue, we design a scalable framework to iterative synthesize data with low cost. Besides, guided tree search algorithms are designed to enable effective “system 2 thinking“ of the prover. HunyuanProver achieves state-of-the-art (SOTA) performances on major benchmarks. Specifically, it achieves a pass of 68.4% on the miniF2F-test compared to 65.9%, the current SOTA results. It proves 4 IMO statements (imo_1960_p2, imo_1962_p2, imo_1964_p2 and imo_1983_p6) in miniF2F-test.
@@ 1 Introduction
We propose HunyuanProver, an LLM-based automatic theorem prover obtained by a pragmatic framework for addressing the aforementioned challenges. The framework takes two core modules: a scalable prover-data generator and guided tree-search algorithms. The prover-data generator only takes open-source data for theorem proving to train the initial autoformalizer and prover. The autoformalizer then converts a large amount of existing math questions into the format of the target prover (e.g., LEAN4). The prover is lastly iteratively improved on such data where new proof data is generated at each iteration to train the prover. At testing time, tree-search algorithms and multiple critic models are designed to perform “slow thinking” that is empirically essential for solving complex theorem-proving tasks.
@@
  • We introduce a scalable pipeline that utilizes open-source data for theorem proving and massive math problems in natural language to generate a large number of new training data for automatic theorem proving.
@@ 4.1 Setup
We use the LEAN engine from LeanDojo (Yang et al., 2024) as the engine to conduct both tactic data generation and benchmark evaluation. Whole and per-step timeout for tactic executing in LEAN are set to 3600 seconds and 60 seconds, respectively. At most 800 search steps are conducted for both BFS and MCTS. For each search step, $K=8$ tactics are sampled from the given LEAN state under temperatures 0.7, 0.8, 1.0 and 1.1, where two tactic are sampled under each temperature. These temperature values are empirically decided. The prompts for policy and distance critic are listed in Appendix B and C.
@@ 4.2 Results and Analysis
Table 1 shows the comparison between HunyuanProver with other existing state-of-the-art systems. Among these systems, our HunyuanProver with distance-tritic as guidance shows the best accuracy on miniF2F, advancing the previous SOTA system (InternLM2.5-StepProver+BFS+CG) by 2.5% points using less search budget. Though InternLM2.5-StepProver+BFS+CG is slightly better than DeepSeek-Prover-V1.5-RL+MCTS, the later does not take any critic for guidance. We believe MCTS can be more beneficial to the task of automatic theorem proving, and this potential has not been fully revealed. In addition, we can also probe the importance of critic guidance by incorporating DCG on HunyuanProver, which shows 3.6%-point gain. We list multiple statements solved by HunyuanProver in Appendix A.